if x+(1/x)=2 then x=? please, answer
Answer by stanbon(75887)

x+(1/x)=2 then x=?
------------
Multiply thru by x to get:
x^2 + 1 = 2x
-----
x^2 - 2x + 1 = 0
-------
Factor:
(x-1)^2 = 0
---
x = 1 with multiplicity 2.
-----------------
Cheers,
Stan H.
=================